In how many ways you can randomly select 5 students from a 12-student class?
Answer by jim_thompson5910(35256)

C(n,r) = (n!)/(r!(n-r)!)

C(12,5) = (12!)/(5!*(12-5)!)

C(12,5) = (12!)/(5!*7!)

C(12,5) = (12*11*10*9*8*7!)/(5!*7!)

C(12,5) = (12*11*10*9*8)/(5!)

C(12,5) = (12*11*10*9*8)/(5*4*3*2*1)

C(12,5) = (95040)/(120)

C(12,5) = 792


There are 792 ways to select 5 students from a 12-student class (assuming order does not matter)